OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = x2 + 8x − 4 (a) (-8,-4) (b) (8, 124) (c) (4, 44) (d) (-4, -20)

OpenStudy (stephstar101):

The lowest point on this graph is at (-4,-20)

OpenStudy (campbell_st):

the line of symmetry passes through the vertex to find it use \[x = -\frac{b}{2a}\] in your question you have \[x = -\frac{8}{2\times1}\] the line of symmetry is x = - 4 substitute x = -4 into the equation to find the y value y = -20 vertex (-4, -20) an alternative method is complete the square y = x^2 + 8x + 16 - 20 y = (x + 4)^2 - 20 vertex at (-4, -20)